OpenStudy (anonymous):

what is the reference angle of -700 degrees what is the reference angle of -700 degrees @Mathematics

jimthompson5910 (jim_thompson5910):

-700 + 2*360 = -700 + 720 = 20 So -700 is coterminal to 20 degrees So the reference angle is 20 degrees